OpenStudy (ksaimouli):

use basic identites and slove sin2 u+tan2 u+cos2 u/sec u

OpenStudy (turingtest):

\[{\sin^2 u+\tan^2 u+\cos^2 u \over \sec u}\]or \[\sin^2 u+\tan^2 u+{\cos^2 u \over\sec u}\]or something else?

OpenStudy (ksaimouli):

the first one

OpenStudy (anonymous):

\[\sin^2(u)+\cos^2(u)=1\] so a start is changing the numerator to \[1+\tan^2(u)\]

OpenStudy (ksaimouli):

sec u

OpenStudy (ksaimouli):

is it sec u

OpenStudy (ksaimouli):

1+tan x/1+cot x can u plz answer this one

OpenStudy (ksaimouli):

ya i got

OpenStudy (turingtest):

what do you have as an answer for the first one?

OpenStudy (ksaimouli):

i got sec u

OpenStudy (turingtest):

good, hmm... let me look at the next one

OpenStudy (turingtest):

ok, \[{1+\tan x \over1+\cot x}\]right?

OpenStudy (ksaimouli):

ya

OpenStudy (turingtest):

multiply top and bottom by tan x, that should do the trick

OpenStudy (ksaimouli):

i did not get u today they said basic identeies so don know any thing about this

OpenStudy (turingtest):

well if you want to talk in terms of identities rewrite it as\[{1+{\sin u \over \cos u}\over1+{\cos u \over \sin u}}\]now multiply top and bottom by\[\tan u={\sin u \over \cos u}\]

OpenStudy (ksaimouli):

ok thx